OpenStudy (anonymous):

The area of a circular sun spot is growing at a rate of 1,500 km^2/s. B) How fast is the radius growing at the instant when the sun spot has an area of 360,000 km^2? HINT [Use the area formula to determine the radius at that instant.] (Round your answer to four decimal places.)

myininaya (myininaya):

So you know the area of circle is?

OpenStudy (anonymous):

A=pi r^2

myininaya (myininaya):

yep and this clearly a rate problem so we will need to find the derivative of that equation with respect to time.

OpenStudy (anonymous):

pi2(r)(dr/dt)

myininaya (myininaya):

and that is what dA/dt is equal to right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so we want to find dr/dt when the area=360000 km^2 and dA/dt=1500 km^2/s

OpenStudy (anonymous):

do we plug it in A'?

myininaya (myininaya):

yep but we still need to know what r is when A=360000

myininaya (myininaya):

Recall A=pi*r^2 so what is r?

myininaya (myininaya):

what is r when A=360000

OpenStudy (anonymous):

.039

myininaya (myininaya):

hmmm I'm not getting that as r

myininaya (myininaya):

\[360000=\pi r^2 \] can you solve this for r?

OpenStudy (anonymous):

600/\[\sqrt{\pi}\]

myininaya (myininaya):

that's good so we have that is r ok and you got earlier that \[A'=2 \pi r r' \]

myininaya (myininaya):

So we have the A' and r now we can find r'

myininaya (myininaya):

replace A' with what it equals replace r with what it equals

OpenStudy (anonymous):

i don't know

myininaya (myininaya):

then solve for r'

myininaya (myininaya):

 what does in those blanks?